Assume that the total score (from both teams) for college football games averages µ = 42 points per game, and that the distribution of total points is approximately normal with σ = 20. What is the likelihood of 4 randomly selected games having an average point total of over 60 points?


1
Expert's answer
2020-10-29T20:04:07-0400

Let "X=" average point total : "X\sim N(\mu,\sigma^2/n)"

Then "Z=\dfrac{X-\mu}{\sigma/\sqrt{n}}\sim N(0,1)"

Given "\mu=42, \sigma=20, n=4"


"P(X>60)=1-P(X\leq60)="

"=1-P(Z\leq\dfrac{60-42}{20/\sqrt{4}})=1-P(Z\leq1.8)\approx"

"\approx1-0.96406968\approx0.03593032"